Could anyone help me with the next questions I sent.
If angle a and angle b is same-side interior angles, and a is 30 degrees, what is angle B?

Found 5 solutions by mananth, MathTherapy, ikleyn, math_tutor2020, n2:
Answer by mananth(16949)
Answer by MathTherapy(10839)


Could anyone help me with the next questions I sent.
If angle a and angle b is same-side interior angles, and a is 30 degrees, what is angle B?
That person is WRONG, as usual! S/he never fails to be WRONG.
The diagram s/he shows is that of INTERIOR-OPPOSITE angles, which indeed, are congruent.
However, same-side INTERIOR angles are supplementary. From this, you should be able to tell ∡B's measure.

Answer by ikleyn(53878)

.
Could anyone help me with the next questions I sent.
If angle a and angle b is same-side interior angles, and a is 30 degrees, what is angle B?
~~~~~~~~~~~~~~~~~~~~~~


        The plot in the post by @mananth is irrelevant to the problem.
        In other words,  it is wrong plot.
        So,  his answer  a = b  is  INCORRECT  and must be denied.


If 'a' and 'b' are same side interior angles, it means that  a + b = 180 degrees.


It means that in this problem 30 + b = 180 degrees.


Hence,  b = 180 - 30 = 150 degrees.    ANSWER

At this point, the problem is solved correctly and completely.



Answer by math_tutor2020(3837)


Same-side interior angles are on the same side of the transversal line, and also inside the parallel likes like so


Since we have parallel lines, the same side interior angles are supplementary. They add to 180 degrees.
a+b = 180
b = 180-a
b = 180-30
b = 150

Tutor mananth would be correct if a,b were alternate interior angles


Answer by n2(89)

.

The same side interior angles, formed by two parallel line and a transversal line,
always sum up to 180 degrees.


So we write

    A + B = 180,

    30 + B = 180,

    B = 180 - 30 = 150 degrees.


ANSWER.  B = 30 degrees.

Solved and explained.

The plot in the post by @mananth is irrelevant to this problem,
so you better ignore it for the safety of your mind.